### v1

# Beyond One World: Benchmarking Super Heros in Role-Playing Across Multiversal Contexts

Abstract
Large language models (LLMs) are increasingly used as role-playing agents, yet their capacity to faithfully and consistently portray version-specific characters—for example, superheroes across comic and cinematic universes—remains underexplored. Superhero canons such as Marvel and DC provide a rich testbed: decades of storytelling yield multiple incarnations of the same character with distinct histories, values, and moral codes. To study this problem, we introduce Beyond One World, a benchmark for character-grounded role-play spanning 30 iconic heroes and 90 canon-specific versions. The benchmark comprises two tasks: (i) Canon Events, which probes factual recall of pivotal life stages, and (ii) Moral Dilemmas, which confronts models with ethically charged scenarios. We score responses for canonical accuracy and reasoning fidelity under a framework that separates internal deliberation (“thinking”) from outward decisions (“acting”). We further propose Think–Act Matching, a metric that quantifies alignment between reasons and actions and serves as a proxy for model trustworthiness. Experiments across reasoning- and non-reasoning-oriented models yield three findings: (1) chain-of-thought prompting improves narrative coherence in weaker models but can reduce canonical accuracy in stronger ones; (2) cross-version generalization within a character remains a major obstacle; and (3) models often excel at either thinking or acting, but rarely both. Beyond One World exposes critical gaps in multiversal consistency and reasoning alignment, offering a challenging evaluation for role-playing LLMs.
11footnotetext: Project repo11footnotetext: Datasets
Beyond One World: Benchmarking Super Heros in Role-Playing Across Multiversal Contexts
Perapard Ngokpol1, Kun Kerdthaisong1, Pasin Buakhaw2,†, Pitikorn Khlaisamniang3, Supasate Vorathammathorn3, Piyalitt Ittichaiwong4,5,*, Nutchanon Yongsatianchot1,* 1Thammasat School of Engineering, Thammasat University 2Department of Computer Engineering and Digital Technology, Faculty of Engineering, Chulalongkorn University 3Artificial Intelligence Association of Thailand 4School of Biomedical Engineering & Imaging Sciences, King’s College London 5Siriraj Informatics and Data Innovation Center (SIData+), Faculty of Medicine, Siriraj Hospital, Mahidol University These authors contributed equally to this work. *Corresponding author
1 Introduction
Large language models (LLMs) have demonstrated strong performance in text generation, translation, and increasingly complex reasoning tasks (Lu et al., 2024). Their growing proficiency has enabled sophisticated applications, including the simulation of personalities and characters (Xi et al., 2023). In character-based role-playing (C-RP), models adopt given personas and are expected to emulate the target’s knowledge, speaking style, and behavior (Chen et al., 2025).
Recent advances in explicit reasoning—such as chain-of-thought (CoT) prompting (Wei et al., 2022) and specialized reasoning models like DeepSeek-R1 (DeepSeek-AI et al., 2025) and Gemini 2.5 Thinking (deepmind, 2025)—further extend LLM capabilities. However, the application of these methods to nuanced role-play remains underexplored (Feng et al., 2025). A central challenge is whether LLMs can consistently portray version-specific characters across contexts (Wang et al., 2024b). Existing benchmarks typically target single-character consistency or basic factual recall. While models can convincingly inhabit a persona in isolated settings, it is unclear whether they maintain coherence as that persona evolves across a narrative timeline—for example, portraying a younger, idealistic incarnation versus a later, more cynical one (Ahn et al., 2024). Moreover, current evaluations rarely analyze the alignment between a model’s reasoning traces and its decisions, across both reasoning-oriented and standard models. Notably, even standard models often produce implicit reasoning (e.g., inner monologue), motivating a more fine-grained assessment of cognitive cues in generated text.
To address these gaps, we focus on superhero characters from the Marvel and DC universes, which offer well-documented, canonically distinct identities spanning timelines and parallel worlds (Stammbach et al., 2022). Decades of storytelling provide detailed histories, evolving moral codes, psychological profiles, motivations, and relationships for numerous versions of the same hero. This rich, curated narrative record yields a complex yet comparatively stable ground truth for evaluating multiversal, version-specific role-play.
In this work, we introduce Beyond One World, a benchmark for character-grounded role-play covering 30 iconic heroes and 90 canon-specific versions. The benchmark assesses two complementary dimensions: (i) Canon Events, which probes factual recall at pivotal life stages (childhood, pre-hero, and established-hero phases) via multiple-choice questions; and (ii) Moral Dilemmas, which presents ethically charged situations inspired by established narrative themes (e.g., save one vs. the greater good, duty vs. personal desire, moral code vs. ends-justify-the-means). We evaluate responses with an LLM-as-a-judge rubric that separates internal deliberation (“thinking”) from outward decisions (“acting”) and score both canonical accuracy and reasoning fidelity.
Our contributions are as follows:
  • Dataset Benchmark: We release Beyond One World, a dataset for version-specific character role-play across 30 heroes and 90 versions, with two tasks—Canon Events and Moral Dilemmas—targeting factual recall and ethically grounded decision-making.
  • Evaluation framework: We propose an analysis framework that disentangles “thinking” from “acting” and applies to both reasoning-oriented and standard LLMs, enabling fine-grained assessment of character-consistent reasoning and decisions.
Dataset/Benchmark Auto-Generated Point-in-time Multiversal Context Reasoning Analysis Items
role-playing same character(movie,series,novel)
SocialBench Chen et al. (2024) 6,493
RoleBench Wang et al. (2024a) 34,523
HPD Chen et al. (2023) 316
CharacterEval Tu et al. (2024) 4,564
TimeChara Ahn et al. (2024) 10,895
Beyond One World(Ours) 2,426
Table 1: Comparison of Benchmarks
2 Related Work
Large language models (LLMs) have recently been used as role-playing language agents (RPLAs) that converse while inhabiting coherent personas. Below, we summarize the main strands of prior work that inform our benchmark.
2.1 From Persona-Chat to Modern RPLAs
Early persona-grounded dialogue framed the task as conditioning a chatbot on a short textual biography. The PersonaChat corpus showed that even simple profiles improve response consistency over generic chit-chat models (Zhang et al., 2018). Subsequent work explored richer conditioning signals—such as goals, memories, and affect—laying the groundwork for contemporary RPLAs.
2.2 Character-Alignment Corpora
More recent datasets target character personas with denser annotations. The Harry Potter Dialogue (HPD) corpus pairs approximately 1k multi-turn scenes with scene metadata, speaker roles, and evolving inter-character relations in English and Chinese, enabling fine-grained character imitation (Chen et al., 2023). In Chinese, CharacterEval provides 1,785 dialogues for 77 literary figures and introduces thirteen metrics spanning conversational ability, consistency, attractiveness, and personality back-testing (Tu et al., 2024).
2.3 Evaluating Sociality and Personalization
Beyond single-speaker fidelity, SocialBench benchmarks 500 personas across roughly 6k scenarios, probing self-awareness, emotion perception, and group adaptability; strong solo performance does not necessarily translate to robust group behavior (Chen et al., 2024). Complementing this, PersoBench measures how well LLMs tailor responses to user-supplied profiles across three dialogue corpora, revealing substantial gaps in personalization despite fluent generation (Afzoon et al., 2024).
2.4 Temporal Consistency on Characters
Maintaining a persona’s limited knowledge at a specific point in time remains challenging. TimeChara evaluates “point-in-time” hallucinations across 10,895 instances and shows that even GPT-4o can leak future facts or inconsistent traits; narrative-expert decomposition mitigates but does not eliminate the issue (Ahn et al., 2024). Our benchmark extends this line by testing multiple universe variants (e.g., childhood, adolescence, pre-hero, hero) of the same character, stressing both temporal and cross-timeline coherence.
3 Dataset Creation
We manually curated a set of well-known hero characters from the Marvel and DC universes, drawing on films, comics, and television series. To avoid role ambiguity, we excluded arcs in which a chosen figure adopts a villainous identity see more details in Appendix E.3. The final collection comprises 30 distinct heroes, each represented in three narrative variants (childhood, pre-hero, and hero phase), yielding 90 character versions in total.
To evaluate an LLM’s factual recall and moral reasoning, we organized the corpus into two tasks: (i) Canonical Events (3.1) probes knowledge of pivotal moments in each hero’s timeline; (ii) Moral Dilemma Situations (3.2) challenge the model to choose actions consistent with the hero’s ethical code;
3.1 Canon Events
This task tests whether an LLM can faithfully recall the key events that define a hero’s backstory. Human annotators with domain expertise wrote four–option multiple-choice questions anchored to pivotal moments in each character’s lore. To reflect a hero’s narrative arc, we split the timeline into three phases: Childhood, spanning early life through adolescence; Pre-Hero, covering the period in which powers are acquired or the first steps toward heroism are taken; and the active Hero phase. For every timeline variant we supply three questions on Childhood, three on Pre-Hero, and nine on the Hero phase, yielding a total of 1,3461{,}346 hand-curated items.
3.2 Dilemma Situations
To gauge whether an LLM can deliberate in line with a hero’s ethical compass, we created a bank of 1,080 multiple-choice dilemmas through a systematic generation and curation process. For each of the 90 character versions, we synthesized three distinct scenarios spanning four archetypal moral conflicts using Claude Sonnet 3.7 to correct grammar, then manually filtered them for canonical fidelity and linguistic clarity.
3.2.1 Dilemma Generation Methodology
Our dilemma generation process employed a structured prompt-based approach designed to ensure both character-specific authenticity and scenario diversity. The generation pipeline consisted of several key components:
Character Integration: Each dilemma was grounded in the specific character’s background, abilities, and established moral framework as derived from their canonical source material. The generation prompt explicitly incorporated the character’s name, lore source, and contextual background to ensure scenario relevance and authenticity.
Iterative Diversification: To prevent redundancy and ensure rich scenario variation, we implemented an iterative generation strategy. For each character-dilemma type combination, three distinct scenarios were created sequentially, with each subsequent iteration explicitly instructed to avoid similarities to previously generated scenarios for that character. This approach yielded scenarios that maintained thematic consistency while exploring different manifestations of the core moral conflict.
Output: Each generated dilemma included:(1) a situational description establishing context and stakes, (2) two binary choice options representing competing moral imperatives, and (3) explicit consequences for each choice. See prompt in (Appendix C.1)
3.2.2 Moral Conflict Taxonomy
We selected four archetypal conflicts that commonly occur in superhero narratives and capture fundamental tensions in moral reasoning across diverse ethical frameworks:
1. Save One vs. Save the Greater Good probes a character’s willingness to sacrifice a few for the many; Iron Man’s self-sacrifice in Avengers: Endgame exemplifies this tension and echoes existential meaning-making through sacrifice (Frankl, 2006) as well as Kohlberg’s post-conventional stage of moral reasoning (Kohlberg, 1971). Generated scenarios in this category typically involved time-sensitive situations where characters must choose between saving a beloved individual versus protecting a larger population, forcing a direct confrontation between particularistic loyalties and universalistic obligations.
2. Hero or Villain captures the struggle between a virtuous self-image and darker impulses—Anakin Skywalker’s fall from grace being the archetype—and draws on Jung’s shadow theory (Jung, 1959), Bandura’s moral disengagement (Bandura, 1999), and Eriksonian identity development (Erikson, 1968). These scenarios explored moments of moral temptation where characters face opportunities to achieve desired outcomes through morally questionable means, testing their commitment to heroic ideals under pressure.
3. Duty vs. Personal Desire mirrors Spider-Man’s dual life and invokes cognitive dissonance (Festinger, 1957), recurring psychosocial role conflict (Erikson, 1968), and Maslow’s hierarchy of needs (Maslow, 1943). Generated dilemmas in this category presented characters with situations where their heroic responsibilities directly conflicted with personal relationships, aspirations, or well-being, examining how they prioritize competing life domains.
4. Ends Justify the Means vs. Moral Code tests whether utilitarian outcomes override deontological constraints; Batman’s refusal to kill despite expedient benefits illustrates the clash between Kantian ethics (Kant, 1785) and the risk of moral injury or ethical fatigue (Litz et al., 2009). These scenarios forced characters to choose between adhering to their established moral principles or compromising them to achieve objectively beneficial outcomes.
Task Datapoints
Canon (270, 270, 806)
Dilemma (270, 270, 270, 270)
Table 2: summary of our benchmark Task 1 consist of 270 Childhood, 270 Pre-Hero and 806 Hero questions(e.g.,Hero1, Hero2, Hero3) while Task 2 consist of balanced Dilemma situations.
Figure 1: Left image is inferencing llm that prompted hero persona to do canon event task, and the right image is prompted llm to do dilemma situation task.
4 Experimental Setups
4.1 Main Experiment
After dataset creation, we designed experiments to evaluate how effectively large language models can adopt character role-playing on hero characters As (Figure 1). Our experimental framework mainly focus on multiple choice question answering with the structured prompting strategies as bellow.
4.1.1 Structured Prompting Design
We developed task-specific prompt templates that provide minimal but essential character context to the models. We employ a consistent base structure.
Initial hero persona You are playing the role of <<name>>, act and think as <<name>>, from <<lore>> <<question>> formatted_question <<question/>>
The <<name>> is character name such Stark “Iron Man” while <<lore>> is source material as the primary identity anchors. lastly formatted_question depends on which task. See Full Task prompt in (Appendix C.2).
4.1.2 Cross-Character Evaluation
To assess the robustness of character differentiation, we implemented a cross-character evaluation protocol. This approach leverages characters that share the same name but originate from different fictional universes (e.g., different movie adaptations, comic series, or TV shows of the same character).
Example
Given a set of characters with identical names but distinct source materials ($CID_{4}$, $CID_{5}$, $CID_{6}$), we systematically evaluate cross-character.
Cross-Character Evaluation Character Set: Spider-Man variants - $CID_{4}$: Spider-Man (Marvel Cinematic Universe, 2016) - $CID_{5}$: Spider-Man (The Amazing Spider-Man 2, 2014) - $CID_{6}$: Spider-Man (Sam Raimi Films, 2002) Evaluation: $CID_{4}$ answers questions originally designed for $CID_{5}$ and $CID_{6}$
4.2 Reasoning analysis framework
We introduce a reasoning analysis framework. Given task outputs from Canon (Section 3.1) and Dilemma (Section 3.2), responses are first segmented into two reasoning traits: <<thinking>> capturing internal deliberation, and <<acting>> capturing decision-making behavior or Physical Appearance. This structured output by Gpt-4o-mini is then judged by Sonnet 3.7, which is prompted to act as a judge scoring.
To scoring the response, prompt conditions the model with character attributes (e.g., Age, Power, Race, Mbti and Enneagram) to simulate a perspective-aligned evaluation process (Appendix C.3). The score value is between 0-5. See the pipeline in (Figure 2)
In addition, to quantify the alignment between a character’s internal reasoning and their outward behavior, we introduce a Think-Act Matching procedure. We embed both <<thinking>> and <<acting>> spans with the all-mpnet-base-v2 Song et al. (2020). Cosine similarity is then computed between the embeddings of each pair. The maximum similarity score across pairs is selected as an indicator of how well a character’s internal reasoning aligns with its external action. This score is interpreted as a proxy for trustworthiness in the character’s response. See result in (Table 3).
Figure 2: A pipeline for reasoning analysis, after getting output from task Canon and Dilemma the output response are structured into <<thinking>> and <<acting>> by Gpt4o-mini judged by Sonnet3.7 that was prompted to be a judge with attributes of that character.
5 Results and Discussion
5.1 Overall Performance on Canon Events
(Table 3) presents results comparing chain-of-thought (CoT) prompting with direct answering (Non-CoT) for the Canon Event task. Several patterns emerge. First, CoT prompting does not consistently improve accuracy: while models such as 4o-mini show a small positive gain (+0.02 accuracy, +0.024 F1), stronger models like sonnet3.5 and sonnet3.7 actually decline in performance when asked to verbalize intermediate reasoning. This suggests that explicit reasoning steps may introduce hallucinations or off-canon elaborations that reduce factual consistency. Interestingly, models such as r1 benefit in terms of reasoning–action alignment (cosine similarity +0.075), even as raw accuracy falls, indicating that CoT can produce more coherent internal deliberation even if the final choice is incorrect.
Model Non-CoT Chain of Thought Difference (CoT - Non-CoT)
Acc F1 Cosim Acc F1 Cosim Acc F1 Cosim
4o-mini 0.626 0.632 0.458 0.646 0.656 0.432 +0.020 +0.024 -0.026
gemini2.5-flash-think 0.663 0.665 0.450 0.388 0.344 0.389 -0.275 -0.321 -0.061
gemini2-flash 0.638 0.647 0.450 0.637 0.652 0.408 -0.001 +0.005 -0.042
r1 0.652 0.656 0.296 0.546 0.573 0.371 -0.106 -0.083 +0.075
sonnet3.5 0.704 0.707 0.433 0.651 0.659 0.404 -0.053 -0.048 -0.029
sonnet3.7 0.637 0.645 0.378 0.620 0.628 0.380 -0.017 -0.017 +0.002
v3 0.615 0.623 0.423 0.502 0.497 0.376 -0.113 -0.126 -0.047
Table 3: Model Performance Comparison: Chain of Thought vs Non-CoT in task canon event. The Cosim score is cosin similarity between matched <<thinking>>-<<acting>> represent trustworthiness of their response.
5.2 Cross-Character Generalization
(Figure 3) shows cross-character transfer. Here, accuracy is substantially lower than within-character evaluations, reflecting the difficulty of distinguishing between overlapping but divergent timelines. Sonnet3.5 again achieves the strongest performance across both Dilemma Cross (0.69) and Canon Cross (0.65), highlighting its robustness to shifts in character context. Conversely, gemini2.5-flash-think shows the steepest drop, indicating vulnerability to timeline conflation. This reinforces that multiversal coherence is not trivially solved by scale or reasoning, but requires models to anchor their decisions in fine-grained contextual cues.
5.3 Thinking vs. Acting
(Figure 4) plots the score that is computed in Section 4.2. Clusters in the upper-right quadrant, showing moderate-to-high fidelity in both dimensions. However, different models emphasize different aspects: gemini2 exhibits stronger reasoning (“thinking” score 3.67) but weaker action alignment, while sonnet3.7 scores highest in acting (3.65) yet shows only moderate internal reasoning (3.03). This divergence illustrates the reasoning–acting gap: models may articulate internally consistent deliberations without translating them into persona-faithful actions, or vice versa.
The Think–Act Matching metric provides further nuance. For example, r1 demonstrates the largest increase in reasoning–action alignment under CoT, despite reduced canonical accuracy. This suggests that reasoning traces can improve trustworthiness of role-played responses, even when they do not strictly improve correctness.
4o-minigemini2.5-flash-thinkgemini2-flashr1sonnet3.5sonnet3.7v300.10.10.20.20.30.30.40.40.50.50.60.60.70.70.80.80.630.630.440.440.550.550.620.620.690.690.600.600.560.560.590.590.420.420.450.450.620.620.650.650.590.590.590.59AccuracyDilemma_Cross and Canon_CrossDilemma CrossCanon Cross
Figure 3: Accuracy result from cross characters evaluation on task canon and task dilemma.
111.51.5222.52.5333.53.544111.51.5222.52.5333.53.544Acting-ScoreThinking-Score4o-minigemini2.5-flash-thinkinggemini2-flashr1sonnet3.5sonnet3.7v3
Figure 4: scores for acting-thinking.
5.4 Discussion
Taken together, these findings highlight several broader insights:
Reasoning CoT helps, but not uniformly: CoT improves coherence for weaker models but can harm stronger ones by over-generating or straying from canon. This echoes (Feng et al., 2025) prior findings on CoT’s mixed effects in knowledge-intensive tasks.
Multiversal consistency is especially hard: Even high-performing models confuse character variants, underscoring the value of our benchmark in testing fine-grained temporal and narrative distinctions.
Acting vs Thinking: A model can stay “in character” in surface-level action without justifying its decisions, or it can reason deeply but act inconsistently. Bridging this gap is key for trustworthy role-play agents.
Overall, our benchmark reveals that current LLMs fall short of fully capturing version-specific character portrayals, especially under timeline shifts and moral dilemmas. These results suggest future work should explore integrated reasoning–persona modeling, potentially combining structured world knowledge with dynamic narrative alignment.
Appendix
Appendix A Our Benchmark
At the end Our Benchmark consist of three tasks: 1.Canon Event(Childhood question for 270, Pre-hero for 270, and hero for 806), 2.Dilemma Situation(“Save one vs. Save the greater good” 270 question, “Hero or Villain” 270 , “Duty vs. Desire 270 ” and “Moral Code vs. Outcome 270 ”).
0200200400400600600800800270270270270806806270270270270270270270270Number of Questions/ItemsBenchmark ItemsCanon EventDilemma Situation
Figure 5: Barplot of our benchmark tasks.The Blue one consist of Childhood,Pre-Hero and Hero question type,the Red one is balanced Dilemma situation.
Appendix B Model Inference Configs
Model Temp OutToken CtxLen
GPT-4o-mini 0.6 1024 128K
Gemini 2.0-Flash 0.6 1024 1M
Gemini-2.5-Flash-Thinking 0.6 1024 1M
Sonnet 3.5 0.6 1024 200k
Sonnet 3.7 0.6 1024 200k
DeepSeek r1 0.6 1024 128k
Sonnet 3.7(judge) 0.1 1024 200k
Table 4: Inference Configs for Answer.
Model Temp OutToken CtxLen
GPT-4o-mini(tags) 0.1 1024 128K
Gemini 2.0-Flash 0.9 1024 1M
Gemini 2.5-Flash-Thinking 0.9 1024 1M
Sonnet 3.7 0.9 1024 200k
Table 5: Inference Configs for Generation.
Appendix C Prompts
C.1 Dilemma Generation
Dilemma Generation You are creating a complex moral dilemma for <<name>> from <<lore>>. The dilemma type is: <<dilemma_type>>. Description of this dilemma type: <<dilemma_descriptions>> Create a binary (two-choice) moral dilemma specific to this character’s background, abilities, and moral framework. The dilemma should force the character to make a difficult choice between two valid but conflicting options. Format your response as a JSON object with these fields: 1. ”situation” – Detailed description of the dilemma scenario (3 sentences) 2. ”choice_A” – First option (1 sentence) 3. ”choice_B” – Second option (1 sentence) 4. ”consequence_A” – Consequence of choosing A (1 sentence) 5. ”consequence_B” – Consequence of choosing B (1 sentence) Make the dilemma deeply personal to this character and challenging based on their specific values and story. Make sure to return a valid JSON object without any additional text before or after.
C.2 Task prompt
Canon Task Prompting:
Canon Task You are playing the role of name, act and think as name, from lore <<question>> canonical_question [option_A], [option_B],[option_C], [option_D] <<question/>>
Dilemma Task Prompting:
Dilemma Task You are playing the role of <<name>>, act and think as <<name>>, from <<lore>>, the situation is [dilemma_scenario] Choice A: [choice_a] Consequence A: [consequence_a] Choice B: [choice_b] Consequence B: [consequence_b] <<question>> [dilemma_type_question] <<question/>>
C.3 Reasoning Analysis
Judge Scoring ”Help me scoring character role-playing, score point between 0-5 , score have two type: thinking(doec thiking response look like reference character) ,acting(does response acting like the character reference), the output must be this format: think_score,act_score example 3,2 Character: <<name>> Source: <<source>> Attributes: <<attributes>> Response to evaluate: <<text_to_process>>
Appendix D Full Results
Table 6, Table 7, Table 9, Table 11
Task Model Acc Prec. Recall F1 Judge_Act Judge_Think Cosim Childhood Pre-Hero Hero1 Hero2 Hero3 canon 4o-mini 0.626 0.646 0.626 0.632 3.191 3.064 0.458 0.581 0.641 0.663 0.593 0.650 gemini2.5-flash-think 0.663 0.682 0.663 0.665 2.972 3.429 0.450 0.700 0.667 0.644 0.689 0.613 gemini2-flash 0.638 0.669 0.638 0.647 3.191 3.673 0.450 0.604 0.648 0.622 0.681 0.635 r1 0.652 0.667 0.652 0.656 3.443 1.789 0.296 0.622 0.733 0.622 0.656 0.628 sonnet3.5 0.704 0.714 0.704 0.707 3.475 3.211 0.433 0.674 0.733 0.700 0.704 0.711 sonnet3.7 0.637 0.671 0.637 0.645 3.652 3.025 0.378 0.630 0.656 0.622 0.656 0.624 v3 0.615 0.648 0.615 0.623 3.250 2.936 0.423 0.607 0.681 0.596 0.604 0.586 canon_cot 4o-mini 0.646 0.680 0.646 0.656 3.183 2.653 0.432 0.626 0.648 0.630 0.637 0.692 gemini2.5-flash-think 0.388 0.463 0.388 0.344 2.578 3.017 0.389 0.352 0.378 0.444 0.381 0.383 gemini2-flash 0.637 0.698 0.637 0.652 2.661 3.443 0.408 0.615 0.663 0.637 0.641 0.632 r1 0.546 0.641 0.546 0.573 2.211 1.423 0.371 0.552 0.585 0.519 0.515 0.560 sonnet3.5 0.651 0.688 0.651 0.659 3.652 3.025 0.404 0.622 0.659 0.630 0.674 0.669 sonnet3.7 0.620 0.656 0.620 0.628 3.405 3.433 0.380 0.607 0.626 0.622 0.630 0.613 v3 0.502 0.569 0.502 0.497 3.250 2.233 0.376 0.544 0.489 0.478 0.500 0.500
Table 6: Canon Event Benchmark Results
Task Model Acc Prec. Recall F1 Judge_Act Judge_Think Cosim Save_Love/Good Hero/Villain Duty/Desire Ends/Code dilemma 4o-mini 0.696 0.731 0.696 0.704 3.302 3.651 0.526 0.548 0.759 0.707 0.770 gemini2.5-flash-think 0.474 0.593 0.474 0.518 2.850 3.837 0.317 0.452 0.433 0.522 0.489 gemini2-flash 0.596 0.663 0.638 0.616 3.446 3.666 0.468 0.481 0.637 0.644 0.622 r1 0.715 0.732 0.715 0.709 3.382 2.405 0.292 0.593 0.778 0.741 0.748 sonnet3.5 0.660 0.686 0.660 0.672 3.723 3.782 0.492 0.574 0.704 0.670 0.693 sonnet3.7 0.571 0.684 0.571 0.618 3.892 4.070 0.400 0.507 0.511 0.648 0.619 v3 0.563 0.602 0.615 0.623 3.638 3.744 0.420 0.552 0.648 0.559 0.493 dilemma_cc 4o-mini 0.669 0.708 0.669 0.684 - - - 0.481 0.778 0.700 0.719 gemini2.5-flash-think 0.459 0.575 0.459 0.505 - - - 0.470 0.430 0.515 0.422 gemini2-flash 0.569 0.664 0.569 0.610 - - - 0.485 0.611 0.485 0.536 r1 0.707 0.720 0.707 0.703 - - - 0.619 0.763 0.730 0.719 sonnet3.5 0.684 0.720 0.684 0.694 - - - 0.604 0.685 0.711 0.737 sonnet3.7 - - - - - - - - - - - v3 0.563 0.608 0.563 0.578 - - - 0.463 0.619 0.567 0.604
Table 7: Dilemma Situations Benchmark Results
Table 8: Accuracy results on canon_cross
Model Accuracy
4o-mini 0.585
gemini2.5-flash-think 0.423
gemini2-flash 0.451
r1 0.619
sonnet3.5 0.650
sonnet3.7 0.591
v3 0.592
Table 9: Accuracy results on canon cross characters evaluation.
Table 10: Accuracy results on dilemma_cross
Model Accuracy
4o-mini 0.634
gemini2.5-flash-think 0.439
gemini2-flash 0.553
r1 0.624
sonnet3.5 0.689
sonnet3.7 0.600
v3 0.558
Table 11: Accuracy results on dilemma cross characters evaluation.
Appendix E Hero Character Profiles
E.1 Hero Selection
In this study, we selected superhero characters from various universes within the Marvel and DC franchises, including the Marvel Cinematic Universe (MCU) Marvel Cinematic Universe Wiki contributors (2025a), the Marvel Animated Universe Marvel Animated Universe Wiki contributors (2025), the Marvel Database Marvel Database Contributors (2025), the DC Extended Universe (DCEU) DC Extended Universe Wiki contributors (2025), and the DC Animated Universe (DCAU) DC Database contributors (2025). The selection process involved filtering for heroes with well-documented backstories who consistently demonstrate pure-hearted intentions and are not categorized as anti-heroes. These criteria were applied to ensure coherence with the scope and objectives of the present research.
E.2 Hero Attributes
During the development of character attributes for the hero evaluation section, we identified five core characteristics: (1) age - the age of the character within the universe up to the latest film or series episode; (2) Power – the hero’s abilities; (3) MBTI – the character’s Myers-Briggs Type Indicator; (4) Race – the character’s species or lineage; and (5) Enneagram – the character’s Enneagram type. General character information, including Age, Power, and Race, was sourced from official and community-curated Wiki fandom pages of each universe. Psychological traits such as MBTI and enneagram types were referenced from PDB: The Personality Database. For example, the character profile of Thor Odinson from the Marvel Cinematic Universe (MCU), shown in Figure 6, lists his age as ’around 1,500 years’, a detail revealed in Avengers: Infinity War and documented on the MCU Wiki (Figure 7). Similarly, Thor’s psychological traits, MBTI and Enneagram, were derived from user-contributed assessments in the Personality Database (Figure 8).
Figure 6: Example of a character profile for Thor Odinson (Thor) from the MCU: The Heroes universe, illustrating the attribute information associated with the character within this universe.
Figure 7: Character age information of Thor Odinson (Thor) from the MCU: The Heroes universe. The age is explicitly referenced in Avengers: Infinity War, where Thor states that he is approximately 1,500 years old. Marvel Cinematic Universe Wiki contributors (2025b)
Figure 8: Character profile webpage of Thor Odin- son (Thor) from the MCU: The Heroes universe from the Personality Database (PDB). This figure shows the online personality profile of Thor Odinson, as portrayed in the Marvel Cinematic Universe (MCU), according to user-contributed data on the Personality Database website. The profile includes MBTI, Enneagram, and other personality traits derived from fan-based assessments. tho
E.3 Question Constructions
To construct a comprehensive benchmark dataset for evaluating.This process involved capturing key aspects of character knowledge and decision-making capabilities in LLM. We developed a custom annotation platform and defined specific structures for leveraging both AI assistance(GPT-4o-mini) and human experts oversight. An annotation platform was developed using Streamlit framework as Figure9. This platform provided a user-friendly interface. It presented expert annotators with structured templates designed to capture information systematically for different benchmarking scenarios. Based on the structured information by the expert Crucially, while GPT-4o-mini assisted in generation, the final validation, refinement, and approval of all data points depends on our expert annotators.
E.3.1 Generation Structure
The annotation process centered around two primary structures.
Canon Event Structure:
This structure format focused on evaluating the model’s ability to recall and accurately represent established facts and events from a character’s history in their lore, by creating multiple-choice Q&A. For each canon event entry, expert annotators were required to complete the following fields within the platform:
  • Character Name and Lore: The name of the hero character and their universe name.
  • Time (When): The specific time context in which the event occurred.
  • Location (Where): The setting or place where the event took place.
  • Description (What happened): A factual account of a significant event from the character’s storyline (canon).
  • Question: A question pertinent to the key event, designed to test the model’s knowledge. This question must be initially drafted by experts.
  • One True Answer: One canonically correct answer by experts and the other answers are generated by GPT-4o-mini.
Dilemma Structure:
This structure aimed to assess the model’s capacity for nuanced role-playing, specifically in navigating complex situations that require decision-making consistent with the character’s established personality, morals, and values. Experts filled out the following fields for each dilemma scenario:
  • Character Name and Lore: The name of the hero character and their universe (which movie or comic).
  • Situation – Time (When): The temporal context for the dilemma.
  • Situation – Location (Where): The setting where the dilemma unfolds.
  • Situation – Context (What is happening): Background information setting the stage and explaining the circumstances leading to the dilemma situation.
  • Dilemma Type: A type of dilemma such as “Save vs. Sacrifice”, “Hero or Villain”, “Duty vs. Desire”.
Figure 9: Platform to help experts construct question and answer to canon event in multiple choices format and construct question in dilemma situation